OpenStudy (anonymous):

The sum of twice a number and another number is 24. The difference of twice the first number and the other number is 12. Which system would model this situation, and what is the solution? A. 2(x + y) = 24 2(x – y) = 12 Solution: (9, 6) B. 2x + y = 24 2x – y = 12 Solution: (9, 6) C. 2(x + y) = 24 2x – y = 12 Solution: (6, 9) D. 2x + y = 24 2(x – y) = 12 Solution: (6, 9)

OpenStudy (anonymous):

is it a ?

OpenStudy (anonymous):

cause i plugged it in and idunno after that.

OpenStudy (shamil98):

The sum of twice a number and another number is 24. The difference of twice the first number and the other number is 12. Which system would model this situation, and what is the solution? 2x + y = 24 2x - y = 12

OpenStudy (shamil98):

yep you got it, and you can give medals to answers btw :D

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

wait nvm lol

OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

tru.

OpenStudy (anonymous):

What is the first step in solving this system of equations by elimination? A. Subtract the two equations. B. Add the two equations. C. Solve for y. D. Multiply each equation by 2.

OpenStudy (shamil98):

what system ? -_-

OpenStudy (anonymous):

2x – 2y = 5 5x + 2y = 9

OpenStudy (shamil98):

B.

OpenStudy (shamil98):

the y cancels each other out and u solve for x..

OpenStudy (anonymous):

ooohhh